a pulse is described by y=4/[(x-3t)^2+8 . what is the velocity of the pulse in si units
Answer #1

Wave equation is hal 2 ozy = ㅈ (A 2x2 zte here speed of wave 4 (8-347-248 +8 a JA oly (1) v=th ya (2-3+)? :.24 = 8CX-3753 = 2